Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
tringali, paoletti, fiori, macchia, soldani - 22:24 Thursday 01 February 2024 (63151) Print this report
Environmenta shift - 25min glitches investigation, HVAC and VAC-UPS switch off

We performed several actions. The list of gps and details are il the attached log file.

switch off of 600N and 600W vacuum pump stations and UPS

we observed a reduction of magnetic noise in CEB, mailing at ~13 Hz sidebads of the 50 Hz and harmonics. Figures 1 and 2.

global site-wide HVAC switch off:

  • NEB and WEB; switched of AHU, cold water pump and chiller
  • CEB: AHU of clean rooms, hall, det lab, inj lab

Figure 3 shows the temperature change and the BNS range. The BNS range progressively improved by ~2Mpc during the switch off (which started at ~16:00 UTC). Yet, this improvement might instead be due to the ongoing thermal transient (TCS RH tuning). The ITF unlocked (IMC unlock) before we could observe the effect of the reverse switch on actions.

disconnection of the cables of the NI and WI ring heaters

Both power supplies were on but set to zero. We disconnected the cables. Once relocked we observed a very large broadband noise in Hrec below ~100 Hz, with a very large 50 Hz. Therefore we connected to ground (tower vac chamber body) all the cables. The broadband noise disappeared, the 50 Hz reduced, but some excess remained. The glitches did not disappeared. We reconnected all cables.

Change the CEB hall temperature

We forced a fast decrease in temperature of the CEB hall by closing the hot water flux and opening to max the cold water flux of the hall HVAC. The temperature reduced by 2 degrees in 40minutes. It seems that the rate of the glitches did not change significantly. To be better analyzed. Figure 4.

 

 

 

 

 

 

 

Images attached to this report
Non-image files attached to this report
Comments to this report:
Paoletti, Fiori, Tringali - 12:11 Friday 02 February 2024 (63153) Print this report
Images attached to this comment
Paoletti, Fiori, Tringali - 15:04 Friday 02 February 2024 (63154) Print this report

About the switch off of 600N and 600W vacuum pump stations and UPS action:

The main magnetic noise coming from the VAC system(s), well visible in ENV_CEB_MAG_V (but also in CEB IPS VOLT and CURR monitors) is a 12.75 Hz broadband noise, and the associated sidebands around the 50 Hz mains (and its harmonics).

A coherence analysis between ENV_CEB_MAG_V and Hrec_hoft does not seem to show significant values (the high value of 25 Hz in Hrec, as well as other peaks, are not associated with the VAC systems but with some action performed on the ITF during the test)

Timing of the operations (UTC):

  • 11:05 600N  pump off
  • 11:10 600W pump off
  • 13:38 600N  rack off
  • 13:53 600W rack off
  • 14:57 600N  UPS off
  • 15:06 600W UPS off
  • 15:10 --> 15:20 quiet time, all VAC 600 susbsistems OFF
  • sequential restart
  • 15:48 all VAC 600 susbsistems ON

Nota bene: looking into the spectrogram it appears that most of the signals that disappear during the shutdown test are peculiar to the 600N station. In fact, the magnetometers are probably close to the IPS cables running to the North Arm.

 

Images attached to this comment
fiori - 18:02 Tuesday 06 February 2024 (63187) Print this report

A look to the noise reduction seen by other magnetometers.

The largest effect was seen by the magnetometer on NI tower base. Instead, the magnetometer at WI tower senses no reduction, and also has overall a quiter magnetic noise than NI.

Images attached to this comment
fiori - 17:51 Wednesday 21 February 2024 (63358) Print this report

To remind that on the day of the HVAC and VAC UPS switch off, magnetometers in CEB were polluted by one extra noise consisting of 1Hz comb noise which was present in the in the CEB UPS currents and was produced by the TCS laser chillers. The contribution of this noise to some magnetic probes can be sized looking at the attached plot. This noise was removed on Feb 20, see https://logbook.virgo-gw.eu/virgo/?r=63357.

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×